THE ELECTRONIC SPECTRUM OF TaCH IN THE VISIBLE REGION

K. ATHANASSENAS, \underline{C. T. KINGSTON}, A. J. MERER AND J. R. D. PEERS, Chemistry Department, University of British Columbia, 2036 Main Mall, Vancouver, BC, Canada V6T 1Z1; M. BARNES, The Fluid Life Corporation, 9619 42nd Avenue, Edmonton, AB, Canada T6E 5R2; D. A. GILLETT, Lambda Physik, Inc., 3201 West Commercial Boulevard, Fort Lauderdale, FL 33309; G. F. METHA, School of Chemistry, University of Sydney, Building F11, Sydney, New South Wales 2006, Australia; L. C. O’BRIEN, Department of Chemistry, Southern Illinois University at Edwardsville, Edwardsville, IL 62026-1652.

Fluorescence excitation spectra of gaseous TaCH and TaCD have been recorded following reaction of laser-ablated tantalum with CH_{4} or CD_{4} under jet-cooled molecular beam conditions. A total of 27 bands of TaCH and 18 of TaCD were observed in the visible region. High resolution spectra reveal the X̃0^{+} ground state to have the substitution structure r_{0}(Ta-C)=1.7714 Åand r_{0}(C-H)=1.080 Å. Since the X̃0^{+} state has no orbital or spin angular momentum, we observe only quadrupole splitting due to the ^{181}Ta nucleus (I=7/2). Resolved fluorescence studies have given the TaCH ground state vibrational frequencies ν_{2}(bend)=640 cm^{−1} and ν_{3}(Ta-C stretch)=960 cm^{−1}. Analysis is currently underway on an Ω = 0^{+} state at 16391.9 cm^{−1} and an Ω = 1 state at 16376.0 cm^{−1}. Through mixing with the Ω = 1 state, the upper Ω = 0^{+} state has acquired some magnetic hyperfine character.